5(x + 4) = 3(8 - x)

Answer by pakhi(24)

5(x + 4) = 3(8 - x)
or 5x + 20 = 24 - 3x
or 5x + 3x = 24 - 20
or 8x = 4
or x = 4/8 = 1/2 = 0.5